# Origin of the Very High Energy -rays in the Low-luminosity Active Galactic Nucleus NGC 4278

### v1

@@ Abstract
NGC 4278, a Low-luminosity active galactic nucleus (AGN), is generally classified as a low-ionization nuclear emission line region (LINER) type AGN. Recently, it is reported to be associated with a very high energy (VHE) -ray source 1LHAASO J1219+2915 in the first Large High Altitude Air Shower Observatory (LHAASO) source catalog. However, no associated counterpart has been detected by Fermi-LAT. By analyzing its X-ray observation data from Swift-XRT, we find it is in a high-flux state on MJD 59546, with the X-ray flux more than one order of magnitude higher than that observed 11.7 year earlier by Chandra. We propose that the detection of VHE -rays from NGC 4278 may be attributed to the presence of an active nucleus displaying behavior similar to that of a BL Lac. To reproduce its spectral energy distribution (SED), we employ a one-zone leptonic model, typically used for fitting broadband SEDs of BL Lacs, and find that smaller values for both Doppler factor () and magnetic field strength () are required than that of typical TeV BL Lacs. Furthermore, NGC 4278 exhibits significantly lower luminosity in both radio and TeV bands when compared with typical TeV BL Lacs. In the radio-luminosity vs. Eddington-ratio plane, NGC 4278 shows greater similarity to Seyfert galaxies and LINERs rather than BL Lacs; however, it still roughly follows the extension towards lower luminosity seen in BL Lacs.
@@ 1 Introduction
Radio-loud active galactic nuclei (RL-AGNs) account for about of the AGN population (Urry & Padovani 1995), characterized as the powerful relativistic jet radiations. Blazars, a special type of RL-AGNs, have their jets pointing nearly along the line of sight and are divided into BL Lacertae objects (BL Lacs) and flat spectrum radio quasars (FSRQs) based on their emission line features in the optical band (Scarpa & Falomo 1997). Blazars are the main extragalactic gigaelectronvolt-teraelectronvolt (GeV–TeV) -ray emitting objects, especially in the TeV band. In recent years, many other types of RL-AGNs have been confirmed as -ray emitters. For example, radio galaxies (Abdollahi et al. 2020), Narrow-line Seyfert 1 galaxies (NLS1s, Abdo et al. 2009; Sun et al. 2015), compact steep-spectrum sources (CSSs, Abdollahi et al. 2020; Zhang et al. 2020), compact symmetric objects (CSOs, Migliori et al. 2016; Principe et al. 2020; Lister et al. 2020; Gan et al. 2021, 2022, 2024), and low-luminosity FR0 type radio galaxies (Grandi et al. 2016; Fu et al. 2022). The large-scale extended regions of jets in AGNs also have been confirmed to emit detectable -rays (Abdo et al. 2010; Ackermann et al. 2016; Yu et al. 2024). It is worth noting that in the TeV band, apart from several FSRQs and radio galaxies, BL Lacs are the main type of detected extragalactic -ray emitting AGNs111http://tevcat2.uchicago.edu/.
@@
Recently, the Large High Altitude Air Shower Observatory (LHAASO) collaboration reported their first catalog (Cao et al. 2024). One new extragalactic TeV -ray emitting source was presented. The point-like source 1LHAASO J1219+2915 located at high Galactic Latitude was detected with a test statistic of TS=49.2 and suggested to be associated with the AGN NGC 4278. The distance between NGC 4278 and 1LHAASO J1219+2915 is 0.05. NGC 4278 is previously classified as a low-ionization nuclear emission line region (LINER, Heckman 1980). No AGNs of this type have been detected in the -ray band.
@@
NGC 4278, located at (González-Martín et al., 2009), belongs to an RL-AGN. However, the total radio luminosity of NGC 4278 is at least 2 orders of magnitude less than the typical RL-AGNs (Condon et al. 1998). The Very Long Baseline Array (VLBA) has resolved its two-sided symmetric steep-spectrum jets on a sub-parsec scale emerging from a flat-spectrum core region (Giroletti et al. 2005). Its radio flux density at 6 cm between 1972 and 2003 exhibits significant variability (100%) on timescale of years, and the outburst around 1985 may be related to the presence of an active nucleus (Giroletti et al. 2005). In the UV band, its nucleus is barely resolved and exhibits evident variability. Its UV flux increases by a factor of 1.6 during a period of 6 months, and the observed UV flare features, amplitude and scale time, are prominently similar to that observed in other low-luminosity AGNs (LLAGNs, Cardullo et al. 2009). Using the Space Telescope Imaging Spectrograph, Cardullo et al. (2009) obtained spectra of the nucleus, measured the broad components of the emission lines, and estimated the mass of the central supermassive black hole (SMBH) to be in the range of . In X-rays, no evident short time-scale (hours and days) variability is observed in the nucleus of NGC 4278 during the Chandra observations. However, an increase in flux by a factor of is observed over a period of 1 hour during the XMM-Newton observation. Younes et al. (2010) also reported that the nucleus flux in NGC 4278 increases by a factor of 3 over a few months and by a factor of 5 between the faintest and brightest observations separated by 3 years. Additionally, NGC 4278 hosts numerous resolved low-mass X-ray binaries (LMXBs), more than one hundred of which exhibit long-term variability (Brassington et al. 2009). In -rays, apart from a transient source, namely 1FLT J1219+2907, reported by the Fermi-LAT transient sources catalog with its position consistent with NGC 4278 from 2009 March 5 to 2009 April 5 (Baldini et al. 2021), no other high-energy -rays have been detected. Therefore, the origin of the detected -ray emission from 1LHAASO J1219+2915 associated with NGC 4278 is confusing.
@@
In this paper, we aim to study the radiation physics and the origin of VHE -rays from NGC 4278 by analysing its observational data in multiwavelength. Data analysis of Fermi-LAT, Chandra, and Swift-XRT is presented in Section 2. In Section 3, we describe the X-ray emission property of the nucleus in NGC 4278. The construction and modeling of the broadband spectral energy distribution (SED) for the nucleus of NGC 4278 are presented in Section 4. Discussion on the origin of VHE -rays and the AGN type of NGC 4278 is given in Section 5, while a summary is provided in Section 6. Throughout, the cosmological parameters km s-1 Mpc-1, , and are adopted.
@@ 2.1 Fermi-LAT
We investigate the latest Fermi-LAT 14 yr Source Catalog (4FGL-DR4, Abdollahi et al. 2022; Ballet et al. 2023) and find only two -ray sources located within a 1 radius circle centered on the radio position of NGC 4278, namely 4FGL J1221.3+3010 and 4FGL J1217.9+3007. They are 0.929 and 0.964 distance from NGC 4278, and are associated with BL Lacs 1ES 1218+304 and B2 1215+30, respectively. Both of these two BL Lacs have been detected in the TeV band and have been confirmed as TeV -ray emitters (Albert et al. 2006; Acciari et al. 2009; Aleksić et al. 2012; Aliu et al. 2013).
@@
To further investigate if there are -ray emission from Fermi-LAT associated with NGC 4278, the Pass 8 data, covering 16 yr from 2008 August 4 to 2024 April 24 (MJD 54682–60424), are extracted from Fermi Science Support Center222https://fermi.gsfc.nasa.gov/ssc/data/access/ for our analysis. We select the region of interest (ROI) centered at the radio position of NGC 4278 (R.A.=185.028, Decl.=29.281) with a radius of 15. The publicly available software Fermitools (version 2.2.0) are used for our analysis. We use event class “SOURCE” (evclass=128) and event type “FRONT+BACK” (evtype=3) for data analysis based on LAT data selection recommendations. The -ray events in the energy range of 0.1–300 GeV are selected with a standard data quality selection criteria of “(DATA_QUAL>0)&&(LAT_CONFIG==1)”. The instrument response function of “P8R3_SOURCE_V3”(Bruel et al. 2018) is used. A maximum zenith angle of is set to reduce the -ray contamination from the Earth limb. The background model in the ROI includes the isotropic background model iso_P8R3_V3_v1.txt and the diffuse Galactic interstellar emission with the parameterized model gll_iem_v07.fits, as well as all the -ray sources listed in the 4FGL-DR4. The normalization of the isotropic background emission and the diffuse Galactic interstellar emission, together with the normalization and spectral parameters of the -ray point sources within a 6.5 radius centered on NGC 4278, are left free, whereas the parameters of those point sources lying beyond are fixed to their 4FGL-DR4 values.
@@
The maximum likelihood test statistic (TS) is used to quantify the significance of the detection of -ray sources, which is defined as (Mattox et al. 1996), where and are the maximum likelihood values for the background with and without a source, respectively. If TS25, it indicates that there is a new source. We generate a residual TS map centered on the radio position of NGC 4278, and find that the maximum TS value in the residual TS map is 13. Therefore, no new -ray source is observed, indicating that no GeV -ray emission from NGC 4278 is detected.
@@ 2.2 Chandra
In this paper, we only analyze the last two observation data from Chandra. The data are reduced with CIAO (v.4.15) and CALDB (v.4.10.4), and then we generate the level-2 event file in a standard procedure. We firstly check the pile-up effect of the two observations by creating pile-up maps with a CIAO tool of pileup_map333https://cxc.harvard.edu/ciao/ahelp/pileup_map.html. No pixel has the count rate per frame greater than 0.1 (corresponding to a pile-up fraction of ), indicating no pile-up effect in the last two observations. We then extract the source photons from a circle centered on the radio position of NGC 4278 with a radius of . The background is taken from an annulus region with inner and outer radii of and , respectively. The spectrum is grouped to have at least 15 counts per bin and the minimization technique is used for the spectral analysis. The spectral fitting is performed using XSPEC (v.12.13.0c). We use the same model adopted in Younes et al. (2010) to fit the spectra in the energy range of 0.5–8.0 keV, i.e., a single power-law absorbed by Galactic and host galaxy plus a thermal component (mekal in XSPEC). The neutral hydrogen column density of Galactic () is fixed as cm-2 (Kalberla et al. 2005) while the value of host galaxy () is set to be free. The fitting parameters of the two spectra are given in Table 1 and the results of the first seven observations reported in Younes et al. (2010) are also presented in the table for comparison. The derived temperatures of the thermal component in the two spectra are almost identical to that reported in Younes et al. (2010). However, we can only obtain an upper limit of for the two observations.
@@ 2.3 Swift-XRT
Swift X-Ray Telescope (Swift-XRT) has performed 4 observations to NGC 4278 in the Photon Counting readout mode between 2017 December and 2021 November. However, the radio position of NGC 4278 is outside of the XRT’s view field during two observations. Therefore, only the other two sets of observational data are used in our analysis; one observation was on 2021 Feb 26 (MJD 59271) with an exposure time of 173 s and another observation was on 2021 Nov 28 (MJD 59546) with an exposure time of 923 s. It is worth noting that the observational data from MJD 59271 have a minimal impact on the spectral fit due to their short exposure time. The result of spectral fit is primarily influenced by the observational data from MJD 59546. The data are processed using the XRTDAS software package (v.3.7.0). The softwave package was developed by the ASI Space Science Data Center and released by the NASA High Energy Astrophysics Science Archive Research Center (HEASARC) in the HEASoft package (v.6.30.1). The calibration files from XRT CALDB (version 20220803) are used within xrtpipeline to calibrate and clean the events. Considering the short exposure time of the two observations, we merge the two XRT event files together using the xselect package and produce a combined average spectrum. The source photons are extracted from a circle centered on the radio position of NCG 4278 with a radius of 20 pixels (). The background is taken from an annulus with an inner and outer radii of 30 pixels () and 45 pixels (), respectively. The combined average spectrum is grouped to have at least 4 counts per bin. We use the public software XSPEC (v.12.13.0c) to fit the spectrum and the C-statistic minimization is adopted for evaluating the goodness of the fits. The energy channels below 0.5 keV and above 8 keV are excluded during fitting. Same as the data analysis of Chandra, the spectrum is fitted by a single power-law absorbed by two absorption components; the neutral hydrogen column density of Galactic is fixed at (Kalberla et al. 2005) while the neutral hydrogen column density of host galaxy () is set to be free. We obtain a C-statistic value/d.o.f of and calculate the corrected flux in the energy range of keV, as listed in Table 1.

### v2

@@ Abstract
NGC 4278, a Low-luminosity active galactic nucleus (AGN), is generally classified as a low-ionization nuclear emission line region (LINER). Recently, it has been reported to be associated with a very high energy -ray source 1LHAASO J1219+2915 in the first Large High Altitude Air Shower Observatory source catalog. However, no associated counterpart has been detected by analyzing the data collected by the Large Area Telescope on board the Fermi Gamma-ray Space Telescope. By analyzing its X-ray observation data from Swift-XRT, we find NGC 4278 is in a high-flux state on MJD 59546, with the X-ray flux more than one order of magnitude higher than that observed 11.7 year earlier by Chandra. Interestingly, this Swift-XRT observation was conducted during the active phase of the -ray source 1LHAASO J1219+2915. We propose that the detection of VHE -rays from NGC 4278 may be attributed to the presence of an active nucleus in its center. To reproduce the spectral energy distribution (SED) of NGC 4278, we employ a one-zone leptonic model, typically used for fitting broadband SEDs of BL Lacs, and find that a smaller magnetic field strength is required than that of typical TeV BL Lacs. Furthermore, NGC 4278 exhibits significantly lower luminosity in both radio and TeV bands when compared with typical TeV BL Lacs. In the radio-luminosity vs. Eddington-ratio plane, NGC 4278 shows greater similarity to Seyfert galaxies and LINERs rather than BL Lacs; however, it still roughly follows the extension towards lower luminosity seen in BL Lacs.
@@ 1 Introduction
Radio-loud active galactic nuclei (RL-AGNs) account for about of the AGN population (Urry & Padovani 1995) and are characterized by powerful relativistic jets. Blazars, a special type of RL-AGNs, have their jets pointing nearly along the line of sight and are divided into BL Lacertae objects (BL Lacs) and flat spectrum radio quasars (FSRQs) based on their emission line features in the optical band (Scarpa & Falomo 1997). Blazars are the main extragalactic gigaelectronvolt-teraelectronvolt (GeV–TeV) -ray emitting objects, especially in the TeV band. In recent years, many other types of RL-AGNs have been confirmed as -ray emitters. For example, radio galaxies (Abdollahi et al. 2020), Narrow-line Seyfert 1 galaxies (NLS1s, Abdo et al. 2009; Sun et al. 2015; D’Ammando 2019; Paliya 2019), compact steep-spectrum sources (CSSs, Abdollahi et al. 2020; Zhang et al. 2020; Gu et al. 2022), compact symmetric objects (CSOs, Migliori et al. 2016; Principe et al. 2020; Lister et al. 2020; Gan et al. 2021, 2022, 2024), and low-luminosity FR0 type radio galaxies (Grandi et al. 2016; Fu et al. 2022). The large-scale extended regions of jets in AGNs also have been confirmed to emit detectable -rays (Abdo et al. 2010a; Ackermann et al. 2016; Yu et al. 2024). It is worth noting that in the TeV band, apart from several FSRQs and radio galaxies, BL Lacs are the main type of detected extragalactic -ray emitting AGNs111http://tevcat2.uchicago.edu/.
@@
Recently, the Large High Altitude Air Shower Observatory (LHAASO) collaboration reported their first catalog (Cao et al. 2024a). One new extragalactic TeV -ray source, 1LHAASO J1219+2915, is included. This source, located at high Galactic latitude, was detected with a significance level of 8.8 during the active period and suggested to be associated with the AGN NGC 4278 (Cao et al. 2024a, b). From a dedicated re-analysis of the LHAASO data, the distance estimated between NGC 4278 and 1LHAASO J1219+2915 is 0.03 (Cao et al. 2024b).
@@
NGC 4278, located at a redshift of (González-Martín et al., 2009), has previously been classified as a low-ionization nuclear emission line region (LINER, Heckman 1980). No AGNs of this type have been detected in the -ray band so far. NGC 4278 belongs to a RL-AGN category. However, the total radio luminosity of NGC 4278 is at least two orders of magnitude lower than that typically observed in RL-AGNs (Condon et al. 1998). The Very Long Baseline Array (VLBA) has resolved its two-sided symmetric steep-spectrum jets on a sub-parsec scale emerging from a flat-spectrum core region (Giroletti et al. 2005). Its radio flux density at 6 cm between 1972 and 2003 exhibits significant variability (100%) on timescale of years, and the outburst around 1985 may be related to the presence of an active nucleus (Giroletti et al. 2005). In the UV band, its nucleus is barely resolved and exhibits evident variability. Its UV flux increases by a factor of 1.6 during a period of 6 months, and the observed UV flare features, amplitude and scale time, are prominently similar to that observed in other low-luminosity AGNs (LLAGNs, Cardullo et al. 2009). Using the Space Telescope Imaging Spectrograph, Cardullo et al. (2009) obtained spectra of the nucleus, measured the broad components of the emission lines, and estimated the mass of the central supermassive black hole (SMBH) to be in the range of . Younes et al. (2010) reported that no evident short time-scale (hours and days) variability is observed in the nucleus of NGC 4278 during the six Chandra observations performed from 2005 February 3 to 2007 April 20. However, an increase in flux by a factor of is observed over a period of 1 hour during the XMM-Newton observation performed on 2004 May 23. Moreover, the nucleus flux in NGC 4278 increases by a factor of 3 over a few months and by a factor of 5 between the faintest and brightest observations separated by 3 years in X-rays. Additionally, NGC 4278 hosts numerous resolved low-mass X-ray binaries (LMXBs), more than one hundred of which exhibit long-term variability (Brassington et al. 2009). In -rays, apart from a transient source, namely 1FLT J1219+2907 (with a photon spectral index of ), reported by the Fermi-LAT transient sources catalog with its position consistent with NGC 4278 from 2009 March 5 to 2009 April 5 (Baldini et al. 2021), no other significant -ray activity has been detected from NGC 4278. Therefore, the origin of the detected -ray emission from 1LHAASO J1219+2915 and its association with NGC 4278 are still debated.
@@
In this paper, we aim to study the radiation physics and the origin of very high energy (VHE) -rays from NGC 4278 by analysing its multiwavelength data. Data analysis of Fermi-LAT, Chandra, and Swift-XRT is presented in Section 2. In Section 3, we describe the X-ray emission property of the nucleus in NGC 4278. The construction and modeling of the broadband spectral energy distribution (SED) for the nucleus of NGC 4278 are presented in Section 4. Discussion on the origin of VHE -rays and the connection to the AGN NGC 4278 is given in Section 5, while a summary is provided in Section 6. Throughout, the cosmological parameters km s-1 Mpc-1, , and are adopted.
@@ 2.1 Fermi-LAT
We investigate the latest Fermi-LAT 14 yr Source Catalog (4FGL-DR4, Abdollahi et al. 2022; Ballet et al. 2023) and find only two -ray sources located within a 1 radius circle centered on the radio position of NGC 4278, namely 4FGL J1221.3+3010 and 4FGL J1217.9+3007. They are 0.929 and 0.964 distance from NGC 4278, and are associated with the BL Lacs 1ES 1218+304 and B2 1215+30, respectively. Both of these two BL Lacs have been detected in the TeV band and have been confirmed as TeV -ray emitters (Albert et al. 2006; Acciari et al. 2009; Aleksić et al. 2012a; Aliu et al. 2013).
@@
To further investigate if there are -ray emission from Fermi-LAT associated with NGC 4278, the Pass 8 data, covering 16 yrs from 2008 August 4 to 2024 April 24 (MJD 54682–60424), are extracted from Fermi Science Support Center222https://fermi.gsfc.nasa.gov/ssc/data/access/ for our analysis. We select the region of interest (ROI) centered at the radio position of NGC 4278 (R.A.=185.028, Decl.=29.281) with a radius of 15. The publicly available software Fermitools (version 2.2.0) and Fermipy (v.1.2.2)(Wood et al., 2017) are used for our binned maximum-likelihood analysis. We use event class “SOURCE” (evclass=128) and event type “FRONT+BACK” (evtype=3) for data analysis based on LAT data selection recommendations333https://fermi.gsfc.nasa.gov/ssc/data/analysis/documentation/Cicerone/Cicerone_Data_Exploration/Data_preparation.html. The -ray events in the energy range of 0.1–300 GeV are selected with a standard data quality selection criteria “(DATA_QUAL>0)&&(LAT_CONFIG==1)”. The instrument response function “P8R3_SOURCE_V3”(Bruel et al. 2018) is used. A maximum zenith angle of is set to reduce the -ray contamination from the Earth limb. The background model in the ROI includes the isotropic background model iso_P8R3_V3_v1.txt and the diffuse Galactic interstellar emission with the parameterized model gll_iem_v07.fits, as well as all the -ray sources listed in the 4FGL-DR4. The normalization of the isotropic background emission and the diffuse Galactic interstellar emission, together with the normalization and spectral parameters of the -ray point sources within a 6.5 radius centered on the radio position of NGC 4278, are left free, whereas the parameters of those point sources lying beyond are fixed to their 4FGL-DR4 values.
@@
The maximum likelihood test statistic (TS) is used to quantify the significance of the detection of -ray sources, which is defined as (Mattox et al. 1996), where and are the maximum likelihood values for the background with and without a source, respectively. We generate a residual TS map centered on the radio position of NGC 4278 and find that the maximum TS value in the residual TS map is 13, while the maximum TS value within a radius around NGC 4278 is approximately 9.
We further extend the analysis by exclusively focusing on PSF3 events above 1 GeV to implement more rigorous event filtering and generate a residual map centered on NGC 4278. The resulting maximum TS value is approximately 4. According to the recommendations from the Fermi collaboration group, LAT catalogs necessitate a minimum TS value of 25 for a source to be deemed significantly detected. Since the significance corresponding to this threshold does not exceed 4, it suggests that the observed signal detection is likely attributable to random fluctuations in the background (Abdo et al. 2010b). The findings suggest that there is no identification of any new -ray source and no detection of significant GeV -ray emission associated with NGC 4278.
In order to investigate the presence of any similar high-energy activity during the active period in the VHE band from 2021 August 23 to 2022 January 10 (MJD 59449–59589, Cao et al. 2024b), we exclusively analyze the LAT data within the energy range of 0.1-300 GeV for this specific time frame. Employing a binned maximum-likelihood analysis approach, we add a point source with a power-law spectral form at the radio position of NGC 4278; however, our findings yield only a TS value of approximately 2.3, indicating no significant detection in the 0.1–300 GeV band during an active period of VHE -ray emission. Therefore, the 2 upper limits of the -ray flux for NGC 4278 in the 0.1–300 GeV band are derived.
@@ 2.2 Chandra
In this paper, we only analyze the last two observation data from the Chandra X-ray Observatory, which are included in the Chandra Data Collection https://doi.org/10.25574/11269 (catalog doi: 10.25574/11269) and https://doi.org/10.25574/12124 (catalog doi: 10.25574/12124), corresponding to obs-ID 11269 on 2010 March 15 and obs-ID 12124 on 2010 March 20. The data are reduced with CIAO (v.4.15) and CALDB (v.4.10.4), and then we generate the level-2 event file in a standard procedure. We firstly check the pile-up effect of the two observations by creating pile-up maps with a CIAO tool of pileup_map444https://cxc.harvard.edu/ciao/ahelp/pileup_map.html. No pixel has the count rate per frame greater than 0.1 (corresponding to a pile-up fraction of ), indicating no pile-up effect in the last two observations. We then extract the source photons from a circle centered on the radio position of NGC 4278 with a radius of . The background is taken from an annulus region with inner and outer radii of and , respectively. The spectrum is grouped to have at least 15 counts per bin and the minimization technique is used for the spectral analysis. The spectral fitting is performed using XSPEC (v.12.13.0c). We use the same model adopted in Younes et al. (2010) to fit the spectra in the energy range of 0.5–8.0 keV, i.e., a single power-law absorbed by Galactic and host galaxy plus a thermal component (mekal in XSPEC). The neutral hydrogen column density of Galactic () is fixed as cm-2 (Kalberla et al. 2005) while the value of host galaxy () is set to be free. The fitting parameters of the two spectra are given in Table 1 and the results of the first seven observations reported in Younes et al. (2010) are also presented in the table for comparison. The derived temperatures of the thermal component in the two spectra are almost identical to that reported in Younes et al. (2010). However, we can only obtain an upper limit of for the two observations.
@@ 2.3 Swift-XRT
Swift X-Ray Telescope (Swift-XRT) has performed 4 observations to NGC 4278 in the Photon Counting readout mode between 2017 December and 2021 November. The analysis is based on only two sets of observational data 555During the other two observations, the radio position of NGC 4278 falls outside the XRT’s view field., specifically obtained on 2021 Feb 26 (MJD 59271) with an exposure time of 173 s and on 2021 Nov 28 (MJD 59546) with an exposure time of 923 s. It is worth noting that the observational data from 2021 Feb 26 (MJD 59271) have a minimal impact on the spectral fit due to the short exposure time. The result of spectral fit is primarily influenced by the observational data from 2021 Nov 28 (MJD 59546). The data are processed using the XRTDAS software package (v.3.7.0). The softwave package was developed by the ASI Space Science Data Center and released by the NASA High Energy Astrophysics Science Archive Research Center (HEASARC) in the HEASoft package (v.6.30.1). The calibration files from XRT CALDB (version 20220803) are used within xrtpipeline to calibrate and clean the events. Considering the short exposure time of the two observations, we merge the two XRT event files together using the xselect package and produce a combined average spectrum. The source photons are extracted from a circle centered on the radio position of NCG 4278 with a radius of 20 pixels (). The background is taken from an annulus with an inner and outer radii of 30 pixels () and 45 pixels (), respectively. We utilize the public software XSPEC (v.12.13.0c) to fit the unbinned spectrum and employ C-statistic minimization to assess the goodness of fit. The energy channels below 0.5 keV and above 8 keV are excluded during fitting to maintain a consistent energy range with that derived from Chandra spectrum fitting. Same as the data analysis of Chandra, the spectrum is fitted by a single power-law absorbed by two absorption components; the neutral hydrogen column density of Galactic is fixed at (Kalberla et al. 2005) while the neutral hydrogen column density of host galaxy is fixed at , which represents an upper-limit value obtained through Chandra spectrum fitting (OBS-ID: 11269). We obtain a C-statistic value/d.o.f of 24.07/30 and calculate the corrected flux in the energy range of keV, as listed in Table 1.